A fish appears to be 2.00 m below the surface of a pond when viewed almost directly above by a fisherman. What is the actual depth of the fish? (n water = 1.33)
Answer #1

The refractive index= real depth/ apparent depth

real depth = refractive index * apparent depth

= 1.33 * 2 = 2.66 m
Answer #2

actual depth of fish is s = nw*s' = 1.33*2 = 2.66m
Answer #3

refractive index= real depth/ apparent depth

=> 1.33 = x/7

=> x = 9.31 m